OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair (a, b). a - b = -5 and a + 2b = 1

OpenStudy (anonymous):

@nansie69x @dtan5457

OpenStudy (dtan5457):

a+2b=1 a-b =-5 Subtract this to eliminate a 3b=6 b=?

OpenStudy (anonymous):

b=3

OpenStudy (dtan5457):

Opps, wrong b=2, not 3

OpenStudy (dtan5457):

a-2=-5

OpenStudy (dtan5457):

a=?

OpenStudy (anonymous):

-3 @dtan5457

OpenStudy (dtan5457):

-3,2 your coordinate

OpenStudy (anonymous):

how about x + y = 16 and x + y = -2?

OpenStudy (dtan5457):

no solution

OpenStudy (anonymous):

2x + 3y = 12 and 3x + 2y = 13? h